The Acme Company lost 650 hours due to accidents on the job in the first quarter of the year. The average hourly wage of the employees who contributed to the lost hours was $15.00. The company benefits add 25% to the wages. Using the cost-of-lost-hours formula (employee hours lost X average loaded labor rate = cost), calculate the direct cost of the lost work hours. Discuss the approach you would use to estimate the hidden costs that would also be associated with the lost hours.
Answer by Theo(13342)

direct labor costs are 15 dollars per hour.

when you add in the cost of benefits, the loaded labor costs are 15 * 1.25 = 18.75 per hour.

assuming that the people were being paid even though they could not work, the company lost 15 * 650 = 9750 in direct labor costs that the workers still receive even though they are not working.

since the company is also paying for benefits, then the company lost 18.75 * 650 = 12187.5 in fully loaded labor costs.

9750 is direct costs paid to the workers.
12187.5 - 9750 = 2437.5 is the cost for additional benefits that the workers still receive even though they are not working.